OpenStudy (anonymous):

A die is rolled twice. Find the probability of rolling a 5 and a 6.

OpenStudy (callisto):

2/6*1/6*2 = 1/9

Directrix (directrix):

Rolling a single die twice is equivalent to rolling a pair of dice once. Looking at the sample space chart a the link above, (5,6) ad (6,5) show among the 36 possibilities. P(5 and 6) = 2/36. I am assuming that the "5" did not have to proceed the "6."